0.1x - 0.09y = 0.18
0.02x + 0.06 y = 0.01

I need to find x y using addition method

Answer by edjones(8007)

0.1x - 0.09y = 0.18
0.02x + 0.06 y = 0.01
.
0.1x-0.09y=0.18
-.1x-.30y =-.05
--------------- Add
====>-.39y=.13
.
3y=1
y=-1/3
x=3/2
.
Ed